Direction: Write an equation for each hyperbola.
vertices: (1,-1) and (1,-9)
conjugate axis of length 6units

I would like to say please and thank you for helping me, in advance.(:

Answer by lwsshak3(11628)

Direction: Write an equation for each hyperbola.
vertices: (1,-1) and (1,-9)
conjugate axis of length 6units
**
This is a hyperbola with vertical transverse axis.
Its standard form of equation: (y-k)^2/a^2-(x-h)^2/b^2=1, (h,k)=(x,y) coordinates of center.
For given hyperbola:
x-coordinate of center=1
y-coordinate of center=(-1+(-9))/2=-10/2=-5 (by midpoint formula)
center: (1,-5)
length of vertical transverse axis=8 (-1 to -9)=2a
a=4
a^2=16
given length of conjugate axis=6=2b
b=3
b^2=9
Equation:
(y+5)^2/16-(x-1)^2/9=1